# RS2V-L: Vehicle-Mounted LiDAR Data Generation from Roadside Sensor Observations

### v1

@@ Abstract
End-to-end autonomous driving solutions, which process multi-modal sensory data to directly generate refined control commands, have become a dominant paradigm in autonomous driving research. However, these approaches predominantly depend on single-vehicle data collection for model training and optimization, resulting in significant challenges such as high data acquisition and annotation costs, the scarcity of critical driving scenarios, and fragmented datasets that impede model generalization. To mitigate these limitations, we introduce RS2V-L, a novel framework for reconstructing and synthesizing vehicle-mounted LiDAR data from roadside sensor observations. Specifically, our method transforms roadside LiDAR point clouds into the vehicle-mounted LiDAR coordinate system by leveraging the target vehicle’s relative pose. Subsequently, high-fidelity vehicle-mounted LiDAR data is synthesized through virtual LiDAR modeling, point cloud classification, and resampling techniques. To the best of our knowledge, this is the first approach to reconstruct vehicle-mounted LiDAR data from roadside sensor inputs. Extensive experimental evaluations demonstrate that incorporating the generated data into model training-complementing the KITTI dataset-enhances 3D object detection accuracy by over 30% while improving the efficiency of end-to-end autonomous driving data generation by more than an order of magnitude. These findings strongly validate the effectiveness of the proposed method and underscore its potential in reducing dependence on costly vehicle-mounted data collection while improving the robustness of autonomous driving models.
@@ III METHOD
This section provides a detailed exposition of the technical implementation of the RS2V-L system. Specifically, Section 3.1 presents the coordinate system transformation and data alignment process, which maps roadside LiDAR data to the vehicle-mounted coordinate frame based on real-time pose estimation. Section 3.2 introduces the methodology for generating virtual LiDAR data at the vehicle end. To achieve accurate scene reconstruction, we employ Patchwork++ [22] for semantic segmentation of object and ground point clouds. The segmented non-ground and ground point clouds are modeled separately, facilitating the synthesis of vehicle-mounted LiDAR point clouds through a virtual radar ray-tracing model. The overall architecture of the proposed method is illustrated in Fig. 2.
Figure 2: The overall technical architecture of RS2V-L.
@@ III-B Vehicle-Mounted LiDAR Data Generation Method
Figure 4: A comparative analysis of the experimental results is presented as follows: (a) illustrates the detection outcomes obtained after training exclusively on the KITTI dataset, where bounding boxes in densely packed vehicle regions exhibit significant clutter and are highly susceptible to misclassification. (b) displays the results following training on both our generated dataset and the KITTI dataset, demonstrating a substantial enhancement in object separation in real-world test scenarios. (c) represents the ground truth annotations. Compared to (a), (b) highlights the effectiveness of our generated dataset in alleviating the challenges associated with dense object occlusion in vehicle-mounted LiDAR datasets, thereby improving detection accuracy and robustness in complex traffic environments.
@@ IV EXPERIMENT
In this study, experiments are conducted using a roadside multi-sensor dataset independently collected by our team. This dataset is acquired through continuous data collection by four 128-line LiDAR sensors deployed at intersections, encompassing a total of 531 frames captured across 50 distinct time periods, including morning and evening rush hours, off-peak hours, and nighttime. Leveraging this dataset, we employ the RS2V-L system to generate 6000 frames of vehicle-mounted point cloud data, covering more than 1000 unique vehicles. On average, each time period captures point cloud data for over 20 different vehicles. The RS2V-L system not only enhances dataset acquisition efficiency by more than an order of magnitude but also significantly enriches data diversity. The generation results are illustrated in Fig. 1.
To assess the effectiveness of the data generated by the RS2V-L method in improving model training, we randomly select 4000 frames for experimentation. The dataset is partitioned into a training set comprising 3000 frames and a test set consisting of 1000 frames. A 3D object detection benchmarking platform is constructed based on OpenPCDet [23], with the KITTI dataset [8] serving as the baseline for comparison. A comprehensive evaluation is conducted from two perspectives: (1) the overall performance enhancement of the model and (2) the specific improvements in complex traffic scenarios, such as densely populated intersections with high traffic flow.
@@ IV-A Overall Model Performance Improvement
To evaluate the impact of the generated data on model performance, a comparative experiment was designed. The control group was trained solely on the KITTI dataset (KITTI: 3712), whereas the experimental group incorporated both the KITTI dataset and an additional 3000 frames of generated data (KITTI: 3712, RS2V-L: 3000) for enhanced training. All training procedures were conducted using the PointPillar [24] model, and vehicle detection performance was assessed on the KITTI validation set. The core evaluation metrics included average precision (AP) and AP_R40 across three key aspects: Bird’s-Eye View (BEV) bounding boxes, 3D bounding boxes, and the target orientation estimation (AOS).
As summarized in TABLE I, the experimental group consistently outperformed the control group across all key metrics. Specifically, the AP values for BEV, 3D, and AOS in the experimental group exhibited a significant increase compared to the control group. Furthermore, improvements were also observed in the AP_R40 metric, indicating that the dataset generated using the RS2V-L method enhances the model’s generalization capabilities across diverse scenarios and driving conditions. These results demonstrate that the RS2V-L framework serves as an effective tool for data augmentation and model optimization, thereby improving the robustness and reliability of autonomous driving perception models.
@@
Indicator Metric Results
KITTI KITTI+RS2V-L
AP BEV 60.9018 78.0783
3D 35.7113 65.4087
AOS 66.89 72.21
AP_R40 BEV 61.4127 80.8127
3D 32.8083 65.3331
AOS 66.89 72.34
@@ IV-B Optimization for Complex Scenarios
To evaluate the effectiveness of the RS2V-L-generated data in enhancing model performance under complex scenarios, as well as to assess its applicability for data collection in extreme conditions, this study integrates the generated dataset into the KITTI training set for joint training. The evaluation focuses on the performance of models tested on the generated vehicle-mounted LiDAR data collected in intersection scenarios. The experimental setup remains consistent with Experiment A, and a comparative study is conducted as follows:
@@
Indicator Metric Results
KITTI KITTI+RS2V-L
AP BEV 1.30 34.98
3D 0.01 28.65
AOS 47.47 44.62
AP_R40 BEV 0.35 31.96
3D 0.01 25.31
AOS 44.74 45.71
According to the test results presented in TABLE II, the average precision (AP) of the 3D detection boxes in the experimental group reaches 28.65, a substantial improvement compared to the mere 0.01 observed in the control group. Similarly, the accuracy of the Bird’s-Eye-View (BEV) detection boxes has also exhibited a significant enhancement. Fig.4 illustrates the visualization results of the experiment, demonstrating that the model trained on the mixed dataset, which incorporates the generated data, achieves a markedly improved object-detection capability. These findings underscore that the RS2V-L-generated data effectively enhances the model’s ability to interpret complex scenarios, thereby improving its performance in object detection under such conditions. The RS2V-L framework proposed in this study leverages 531 frames of roadside LiDAR data to generate 6000 frames of vehicle-mounted LiDAR data, achieving a more than tenfold increase in data collection efficiency. Furthermore, the effectiveness of the RS2V-L-generated data in improving model performance is systematically validated through two sets of comparative experiments.
In the overall performance evaluation, the model trained with the augmented dataset, which incorporates the generated data, consistently outperforms the model trained solely on the KITTI dataset across multiple key metrics, including Bird’s-Eye-View (BEV), 3D detection, and the Angle of Object Sighting (AOS). Specifically, the BEV metric improves from 60.91 to 78.09, reflecting a 28-percentage-point increase; the 3D detection accuracy rises from 35.71 to 65.41, marking an 83-percentage-point enhancement; and the AOS metric increases from 66.89 to 72.71, representing an 8-percentage-point gain. In the complex scenario evaluation, particularly in intersection environments, the model trained with the augmented dataset significantly surpasses the baseline model trained solely on KITTI data in both 3D detection and BEV accuracy, demonstrating its robustness in handling challenging traffic conditions. In summary, the RS2V-L-generated data not only enhances the overall model performance but also substantially improves its effectiveness in complex and highly dynamic scenarios. This approach offers significant value for augmenting vehicle-mounted LiDAR datasets and facilitating data collection in extreme driving conditions.
@@ V CONCLUSIONS
This paper presents the RS2V-L method, which effectively reconstructs and synthesizes vehicle-mounted LiDAR data from roadside sensor observations, offering an efficient and cost-effective data generation paradigm for autonomous driving applications. Experimental evaluations demonstrate that the data generated by RS2V-L not only significantly enhances the overall performance of autonomous driving models but also exhibits remarkable adaptability in complex environments, such as intersections.
In contrast to conventional vehicle-mounted data acquisition approaches, RS2V-L capitalizes on the "god’s eye view" advantage of roadside sensors to mitigate vehicle-to-vehicle occlusion while addressing domain adaptation challenges between simulated and real-world data. Furthermore, this method accommodates arbitrary vehicle models and sensor configurations, ensuring broad applicability across diverse autonomous driving scenarios.
Future research will focus on further optimizing the RS2V-L framework, expanding its applicability to a wider range of scenarios and sensor modalities, and integrating multimodal data generation techniques. These advancements aim to provide more comprehensive support for the continued development of autonomous driving technology.

### v3

# RS2AD: End-to-End Autonomous Driving Data Generation from Roadside Sensor Observations

@@ Abstract
End-to-end autonomous driving solutions, which process multi-modal sensory data to directly generate refined control commands, have become a dominant paradigm in autonomous driving research. However, these approaches predominantly depend on single-vehicle data collection for model training and optimization, resulting in significant challenges such as high data acquisition and annotation costs, the scarcity of critical driving scenarios, and fragmented datasets that impede model generalization. To mitigate these limitations, we introduce RS2AD, a novel framework for reconstructing and synthesizing vehicle-mounted LiDAR data from roadside sensor observations. Specifically, our method transforms roadside LiDAR point clouds into the vehicle-mounted LiDAR coordinate system by leveraging the target vehicle’s relative pose. Subsequently, high-fidelity vehicle-mounted LiDAR data is synthesized through virtual LiDAR modeling, point cloud classification, and resampling techniques. To the best of our knowledge, this is the first approach to reconstruct vehicle-mounted LiDAR data from roadside sensor inputs. Extensive experimental evaluations demonstrate that incorporating the data generated by the RS2AD method (the RS2V-L dataset) into model training as a supplement to the KITTI dataset can significantly enhance the accuracy of 3D object detection and greatly improve the efficiency of end-to-end autonomous driving data generation. These findings strongly validate the effectiveness of the proposed method and underscore its potential in reducing dependence on costly vehicle-mounted data collection while improving the robustness of autonomous driving models.
@@ III METHOD
This section provides a detailed exposition of the technical implementation of the RS2AD system. Specifically, Section 3.1 presents the coordinate system transformation and data alignment process, which maps roadside LiDAR data to the vehicle-mounted coordinate frame based on real-time pose estimation. Section 3.2 introduces the methodology for generating virtual LiDAR data at the vehicle end. To achieve accurate scene reconstruction, we employ Patchwork++ [22] for semantic segmentation of object and ground point clouds. The segmented non-ground and ground point clouds are modeled separately, facilitating the synthesis of vehicle-mounted LiDAR point clouds through a virtual radar ray-tracing model. The overall architecture of the proposed method is illustrated in Fig. 2.
Figure 2: The overall technical architecture of RS2AD.
@@ III-B Vehicle-Mounted LiDAR Data Generation Method
Figure 4: Two scene examples are depicted, with the ground - truth illustrated on the far left. For each example, the first row presents the test results of distinct models (PointPillars, SECOND, PV-RCNN) trained solely on the KITTI dataset, while the SECOND row showcases the test outcomes of these models when trained on a combination of the KITTI dataset and the RS2V-L dataset. Through comparison between the two rows, a notable enhancement in detection accuracy is observable for the models trained with the mixed - dataset strategy.
@@ IV EXPERIMENT
In this study, experiments are conducted using a roadside multi-sensor dataset independently collected by our team. This dataset is acquired through continuous data collection by four 128-line LiDAR sensors deployed at intersections, encompassing a total of 531 frames captured across 50 distinct time periods, including morning and evening rush hours, off-peak hours, and nighttime. Leveraging this dataset, we employ the RS2AD system to generate 6000 frames of vehicle-mounted point cloud data, covering more than 1000 unique vehicles. On average, each time period captures point cloud data for over 20 different vehicles. The RS2AD system not only enhances dataset acquisition efficiency by more than an order of magnitude but also significantly enriches data diversity. The generation results are illustrated in Fig. 1.
To assess the effectiveness of the data generated by the RS2AD method in improving model training, we randomly select 4000 frames for experimentation. The dataset is partitioned into a training set comprising 3000 frames and a test set consisting of 1000 frames. A 3D object detection benchmarking platform is constructed based on OpenPCDet [23], with the KITTI dataset [8] serving as the baseline for comparison. A comprehensive evaluation is conducted from two perspectives: (1) the overall performance enhancement of the model and (2) the specific improvements in complex traffic scenarios, such as densely populated intersections with high traffic flow.
@@ IV-A Overall Model Performance Improvement
To evaluate the impact of the generated data on model performance, a comparative experiment was designed for models, including PointPillars[24], SECOND, and PV-RCNN. Each model was trained under three configurations: solely on the KITTI dataset (KITTI: 3712), solely on the RS2V-L dataset, or on a mix of KITTI (3712 frames) and RS2V-L (3000 frames). Vehicle detection performance was evaluated on the KITTI validation set. The core evaluation metrics comprised average precision (AP) and AP_R40 for Bird’s-Eye View (BEV) bounding boxes and 3D bounding boxes under Intersection over Union (IoU) thresholds of 0.7 and 0.5.
As summarized in TABLE I, for all models, the mix-training group (KITTI + RS2V-L) exhibited significant performance improvements. For instance, when compared to training with the KITTI dataset alone, the 3D bounding-box AP of the SECOND model under an IoU of 0.7 increased from 65.08 to 67.29 in the mix group. Compared to training with the RS2V-L dataset alone, the 3D bounding-box AP of PointPillars under an IoU of 0.5 rose from 58.61 to 84.32 in the mix group. These improvements in the AP and AP_R40 metrics illustrate that mixed training with RS2V-L enhances model performance and generalization. Thus, the RS2AD framework serves as an effective data-augmentation and model-optimization tool, enhancing the robustness and reliability of autonomous-driving perception models.
@@
Model Dataset IoU: 0.7 IoU: 0.5
AP AP_R40 AP AP_R40
bev 3d bev 3d bev 3d bev 3d
PointPillars Kitti 79.27 65.87 79.83 67.44 84.65 84.29 87.93 86.48
RS2V-L 54.21 42.22 54.33 41.16 59.72 58.61 60.27 59.27
Mix 79.19 65.58 79.70 66.90 84.80 84.32 88.10 86.48
SECOND Kitti 77.74 65.08 78.80 66.25 83.42 82.88 86.28 84.73
RS2V-L 59.34 52.59 60.50 52.94 63.64 63.19 64.94 64.34
Mix 77.43 67.29 78.95 67.62 83.59 83.21 86.50 84.97
PV-RCNN Kitti 79.52 70.92 80.31 71.26 83.7 83.60 87.24 85.77
RS2V-L 27.59 3.88 26.60 3.19 55.00 50.14 52.56 47.62
Mix 77.66 71.50 78.78 71.36 81.90 81.81 85.19 84.65
@@ IV-B Optimization for Complex Scenarios
To evaluate the effectiveness of the RS2AD-generated data in enhancing model performance under complex scenarios and its applicability for data collection in extreme conditions, this study conducts experiments with models such as PointPillars, SECOND, and PV-RCNN. Each model is trained under three configurations: solely on the KITTI dataset, solely on the RS2V-L dataset, or on a mix of KITTI and RS2V-L. The evaluation focuses on model performance when tested on the RS2V-L test set.
As shown in TABLE II, for all models, mixed training (KITTI + RS2V-L) yields significant performance improvements. Take PointPillars as an example: compared to KITTI only training, its 3D detection-box average precision (AP) on the RS2V-L test set improves remarkably; compared to RS2V-L only training, the mix-training group also demonstrates notable enhancements. These AP metric improvements for 3D and Bird’s-Eye-View (BEV) detection boxes illustrate that incorporating RS2V-L for mixed training effectively strengthens the model’s capability to handle complex scenarios, thereby enhancing object-detection performance. Fig.4 displays the experiment’s visualization results, where the model trained on the mixed dataset shows a significantly improved object-detection ability, further confirming the positive impact of RS2AD-generated data.
@@
Model Dataset IoU: 0.7 IoU: 0.5
AP AP_R40 AP AP_R40
bev 3d bev 3d bev 3d bev 3d
PointPillars Kitti 14.31 9.31 8.76 0.81 34.09 28.02 31.16 24.54
RS2V-L 77.09 61.58 77.47 60.55 88.44 86.19 89.70 88.68
Mix 77.23 60.94 77.17 66.45 87.94 86.23 90.13 88.08
SECOND Kitti 13.08 4.65 6.49 0.97 28.66 23.87 24.98 19.85
RS2V-L 77.82 64.24 79.39 63.10 88.41 87.52 91.13 88.70
Mix 78.01 61.21 79.36 60.58 88.29 86.69 90.56 88.47
PV-RCNN Kitti 17.17 9.56 11.88 2.69 32.01 27.58 28.89 24.12
RS2V-L 73.57 60.51 75.57 60.01 85.07 83.52 87.14 85.07
Mix 78.22 66.15 79.70 65.07 88.01 86.68 89.28 88.73
The RS2AD framework proposed in this study leverages 531 frames of roadside LiDAR data to generate 6000 frames of vehicle-mounted LiDAR data, achieving a more than tenfold increase in data collection efficiency. Furthermore, the effectiveness of the RS2AD-generated data in improving model performance is systematically validated through two sets of comparative experiments.
In the overall performance evaluation, models such as PointPillars, SECOND, and PV-RCNN trained on the augmented dataset (KITTI + RS2V-L) overall outperform their counterparts trained solely on the KITTI dataset, across key metrics for both Bird’s-Eye-View (BEV) and 3D object detection under various Intersection over Union (IoU) thresholds. Taking SECOND as an illustrative example, under an IoU threshold of 0.7, the 3D average precision (AP) improves from 65.08 (KITTI-only) to 67.29 (Mix). At a more relaxed threshold of IoU = 0.5, the 3D AP rises from 83.42 to 83.59.
In the evaluation under complex scenarios—particularly at intersection-like environments assessed on the RS2V-L test set—the benefits of the augmented dataset are even more pronounced. Specifically, for PV-RCNN under IoU = 0.7, the BEV AP improves dramatically from 17.17 (KITTI only) to 78.22 (Mix), and the 3D AP increases from 9.56 to 66.15. Under IoU = 0.5, the BEV AP rises from 32.01 to 88.01, while the 3D AP climbs from 27.58 to 86.68.
In summary, the RS2AD-generated data not only enhances the overall model performance but also substantially improves its effectiveness in complex and highly dynamic scenarios. This approach offers significant value for augmenting vehicle-mounted LiDAR datasets and facilitating data collection in extreme driving conditions.
@@ V CONCLUSIONS
This paper presents the RS2AD method, which effectively reconstructs and synthesizes vehicle-mounted LiDAR data from roadside sensor observations, offering an efficient and cost-effective data generation paradigm for autonomous driving applications. Experimental evaluations demonstrate that the data generated by RS2AD not only significantly enhances the overall performance of autonomous driving models but also exhibits remarkable adaptability in complex environments, such as intersections.
In contrast to conventional vehicle-mounted data acquisition approaches, RS2AD capitalizes on the "god’s eye view" advantage of roadside sensors to mitigate vehicle-to-vehicle occlusion while addressing domain adaptation challenges between simulated and real-world data. Furthermore, this method accommodates arbitrary vehicle models and sensor configurations, ensuring broad applicability across diverse autonomous driving scenarios.
Future research will focus on further optimizing the RS2AD framework, expanding its applicability to a wider range of scenarios and sensor modalities, and integrating multimodal data generation techniques. These advancements aim to provide more comprehensive support for the continued development of autonomous driving technology.